OpenStudy (anonymous):

The equation of a line is y=-1/2x-2. What is the equation of the line that is perpendicular to the first line and passes through the point (2, –2)? A. y = 2x – 2 B. y=1/2x - 3 C. y = 2x – 6 D. y = –2x + 2

OpenStudy (anonymous):

The slope in the equation given is -1/2. Since we are looking for a perpendicular line, we need he negative reciprocal of that slope. All that means is " flip " the slope and change the sign. Therefore, the negative reciprocal of -1/2 is 2/1 or just 2. now we use y = mx + b (and solve for b) slope(m) = 2 (2,-2) x = 2 and y = -2 now we sub y = mx + b -2 = 2(2) + b -2 = 4 + b -2-4 = b -6 = b so your perpendicular line is : y = 2x - 6

OpenStudy (anonymous):

Thank you!